Suppose mass (m = 482 g) hanging in the middle divides the string (length L = 46.1 cm) exactly in half (and the tension in the two strings are the same), and the distance between the ends of the string (on top) is 25.7 cm. a) Find the magnitude of the tension in either string.
Answer #1

d = 25.7 cm , x = d/2, L = 46.1 cm

cos(theta) = x/L = 25.7/(2*46.1)

From Newton second law net force along verticla direction is zero

2Tcos(theta) = mg

2*T*25.7/(2*46.1) =0.482*9.8

T =8.47 N